The Frush Corporation has two different bonds currently outstanding. Bond M has a face value of $60,000 and matures in 20 years. The bond makes no payments for the first six years, then pays $2,700 every six months over the subsequent eight years, and finally pays $3,000 every six months over the last six years. Bond N also has a face value of $60,000 and a maturity of 20 years; it makes no coupon payments over the life of the bond. The required return on both these bonds is 12 percent compounded semiannually.

  

What is the current price of Bond M and Bond N? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Current price
  Bond M $   
  Bond N $   
Answer #1

Current price of Bond M=2700/1.06^13*(1-1/1.06^16)/(1-1/1.06)+3000/1.06^29*(1-1/1.06^12)/(1-1/1.06)+60000/1.06^40=$24,313.99

Current price of Bond N=60000/1.06^40=$5,833.33